Calculate the net present value of the following project for discount rates of 0, 50, and 100%: (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations. Round your answers to 2 decimal places.)

C0 C1 C2
−$7,233.75 +$4,680.00 +$19,575.00
Discount Rate NPV         
0%           $   
50%             
100%             
b. What is the IRR of the project?
  IRR %  
Answer #1

NPV is given by:

Answer a)

Discount rate = 0%

NPV = [ 4680 / (1 + 0%)^1 ] + [ 19575 / (1 + 0%)^2 ] - Initial Investment

NPV = 4680 + 19575 - 7233.75

NPV = $17021.25

Discount rate = 50%

NPV = [ 4680 / (1 + 50%)^1 ] + [ 19575 / (1 + 50%)^2 ] - Initial Investment

NPV = 3120 + 8700 - 7233.75

NPV = $4586.25

Discount rate = 100%

NPV = [ 4680 / (1 + 100%)^1 ] + [ 19575 / (1 + 100%)^2 ] - Initial Investment

NPV = 2340 + 4893.75 - 7233.75

NPV = $0

Discount Rate (r) NPV
0% 17021.25
50% 4586.25
100% 0

-------------------------------------------------

Answer b)

In answer (a), we found that at discount rate 100%, we have NPV = 0

Hence, IRR = 100%

At irr

NPV = 0

NPV = 0 = [ 4680 / (1 + r%)^1 ] + [ 19575 / (1 + r%)^2 ] - Initial Investment

0 = [ 4680 / (1 + 100%)^1 ] + [ 19575 / (1 + 100%)^2 ] - 7233.75

Solving for r

r = 100%